The length of a rectangle is 10 in. more than twice its width. If the perimeter of the rectangle is 38 in., find the width of the rectangle
Answer by scott8148(6628)

let x=width, so 2x+10=length __ 2(x)+2(2x+10)=38 __ 2x+4x+20=38 __ 6x=18